OpenStudy (anonymous):

, what are the extended values of the function? y=4/9x-45

OpenStudy (anonymous):

9(y+45)=4/x y=-405x-4/9x 9x(y+45)=4 solution is x=0 y=4-405x/9x

OpenStudy (anonymous):

My answer choices are -5, 0.2, 5, -0.2.

OpenStudy (anonymous):

Is the denominator 5x-45?